a coop wants to use 192 feet of fencing to put a fence around a rectangular garden. the garden is going to be twice as wide as it is long. what are the dimensions of the rectangular garden?
Answer by macston(5194)

L=Length; W=Width=2L; P=perimeter=192 feet
P=2(L+W) Substitute for W.
192 feet=2(L+2L)
192 feet=2(3L)
192 feet=6L Divide each side by 6.
32 feet=L ANSWER 1: The length is 32 feet
W=2L=2(32 feet)=64 feet ANSWER 2: The width is 64 feet.
CHECK:
P=2(L+W)
192ft=2(32ft+64ft)
192ft=2(96ft)
192ft=192ft